OpenStudy (anonymous):

Express the exact value of csc 60°, with a rational denominator.

OpenStudy (anonymous):

cosecx = 1/sinx cosec60 = 1/sin60 As sin60 = root3/2 So, cosec60 =\[1/\sqrt{3}/2\] \[cosec60 = 2/\sqrt{3}\] Rationalising we get, \[cosec60 = 2\sqrt{3}/3\] is your answer..

OpenStudy (anonymous):

thanks!

OpenStudy (anonymous):

Welcome dear...